# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v1

@@ Appendix A SUPPLEMENTAL MATERIAL
Appendix B Cluster density profiles and constraint
Figure 5: Left: Density profiles of cluster host halos in the CDM, SIDM2v, and SIDMx simulations (dashed). All profiles are well fitted by a two-component parametric SIDM model (solid) Hou et al. (2025); Yang et al. (2024). Right: Evolution of an isolated cluster host halo, taking scale parameters from the cluster halo in the CDM simulation, under SIDM2v. At either $t=5~{}\rm Gyr$ (solid) or $14~{}\rm Gyr$ (dotted), the resulting core (blue) is much smaller than that produced in a one-component model with $\sigma/m=0.1~{}\rm cm^{2}/g$ (SIDM01, red), demonstrating that SIDM2v remains consistent with cluster-scale constraint.
@@ Appendix B Cluster density profiles and constraint
To demonstrate the interplay between inter- and intra-species interactions, and to show that cluster halos in SIDM2v remain consistent with the current $0.1~{}\rm cm^{2}/g$ constraint, we present the density profiles of host halos from cosmological simulations in the left panel of Fig. 5. We also simulate an isolated cluster halo under SIDM2v, using the scale parameters from the CDM cluster halo as a reference. The resulting density profiles are denser than those of the one-component SIDM01 model at both $t=5$ and $14~{}\rm Gyr$. Since the observational constraint is based on the inner mass distribution, our results indicate that clusters in SIDM2v are less constrained than in SIDM01.
Appendix C Calculation of strong lensing effects
Figure 6: Density profiles of all cluster subhalos with $M>10^{11}\,M_{\odot}/h$ in the CDM, SIDM2v, and SIDMx simulations (dashed). Solid curves show corresponding fits using the parametric SIDM model of Refs. Hou et al. (2025); Yang et al. (2024). Subpanels display the relative deviation between simulation and fit, defined as $2\,(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The parametric model accurately captures the density structure of simulated subhalos across a broad mass range.
We compute strong lensing effects using the parametric approach developed in Refs. Hou et al. (2025); Yang et al. (2024). Starting from the simulated particle data, we extract spherically averaged halo density profiles under the assumption of spherical symmetry. These profiles are then modeled using a sum of two parametric SIDM (PSIDM) halo components, each described by an analytic fitting function. For each PSIDM component, the density profile is given by
$\rho_{\rm PSIDM}(r|\rho_{s,0},r_{s,0},\tau)$ (1)
where $\rho_{s}=\rho_{s}(\rho_{s,0},\tau)$, $r_{s}=r_{s}(r_{s,0},\tau)$, $r_{c}=r_{c}(r_{s,0},\tau)$, $\beta=\beta(\tau)$ and $\gamma=\gamma(\tau)$ are functions of the phase parameter $\tau\in(0,1.08)$, defined explicitly in Ref. Hou et al. (2025). Since each PSIDM component is determined by three parameters, the sum of two such components involves six fitting parameters in total.
Figure 6 shows the fitting performance for cluster subhalos in three cosmological simulations: from left to right, CDM (blue), SIDM2v (red), and SIDMx (magenta). The lower panels illustrate the relative difference between the simulated (Sim) and fitted (Fit) profiles, computed as $2(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The two-component PSIDM model reproduces the simulated profiles well, and the remaining scatter is primarily due to statistical fluctuations in the simulation data.
Based on the fitted profiles, Ref. Hou et al. (2025) provides analytic expressions for the surface density, lensing potential, and deflection angles of each halo. The linearity of the lensing equation allows us to stack their contributions for more comprehensive lensing predictions.
For halos other than the main cluster hosts, their internal ellipticities contribute subdominantly to the lensing shear compared to the external shear from the host halo. However, the mass and ellipticity of the host halo significantly affect the lensing signal and must be incorporated. We model this effect by applying a coordinate transformation that introduces a fixed ellipticity into the projected mass distribution. The ellipticity of each halo is measured by projecting the associated particles onto the lens plane, taken to be the $x$$y$ plane of the simulation box. We define the axis ratio $q=b/a$ of the projected mass distribution and compute the elliptical radial coordinate as Kormann et al. (1994)
$s=\sqrt{q^{2}(x\cos\phi+y\sin\phi)^{2}+(-x\sin\phi+y\cos\phi)^{2}},$
where $\phi$ is the angle between the $x$-axis and the semi-major axis of the lens. Substituting $s$ for $r$ in the circular lens model yields an elliptical lens, preserving the enclosed mass within isodensity contours. We apply a fast Fourier transform to the elliptical surface density to compute the deflection angle maps for the host halos. For the CDM, SIDM2v, and SIDMx clusters, the measured axis ratios are $q=0.39$, $0.42$, and $0.47$, respectively. These large ellipticities enhance the GGSL effects associated with the host halos.
Appendix D Strong lensing effects under identical host lens
Figure 7: Left: Tangential (opaque) and radial (40% transparent) critical curves (red) overlaid on projected density maps in the SIDM2v (top) and SIDMx (bottom) simulations, where the host halo has been replaced with the CDM host halo to isolate subhalo contributions. Right: Tangential caustics (green) correspond to the critical curves on the left panels. The enhancement in the number and strength of small-scale lenses remains clearly demonstrated, confirming that subhalo structure is the primary driver of GGSL cross section differences between CDM and two-component SIDM models. See text for details.
To address the excess in the GGSL cross section, we primarily focus on the influence of cluster subhalos. To assess the impact of uncertainties in modeling host halos, we perform an extreme test by adopting the CDM cluster as the host lens in all cases. This approach helps to marginalize the contribution from host halo variations.
Figure 7 shows the critical curves (left) and caustics (right) for the SIDM2v and SIDMx simulations, with their host halos replaced by the CDM host. The presence of a significant number of small-scale lenses remains visually apparent through the enclosed curves. In the SIDM2v case, more structures are visible in the upper-left region, which can be attributed to a broader spatial distribution of massive subhalos compared to SIDMx. We leave a detailed investigation of the origin of these differences, likely related to variations in accretion history, for future work with higher-resolution simulations.
Under this setup, the resulting GGSL cross sections are $0.515$, $11.7$, and $6.50$ for the CDM, SIDM2v, and SIDMx simulations, respectively. The corresponding numbers of secondary lenses are $5$, $25$, and $26$. Comparing these results with those presented in the main text, we find that the choice of cluster halo can indeed introduce substantial differences. The differences in the simulated host halos can be physical. However, future simulations with higher resolution and full baryonic physics may further refine these details. Despite this, our conclusions remain robust: the ratio of $\sigma_{\rm GGSL}$ between SIDM2v and CDM increases from $13$ to $23$. The number of small-scale lenses decreases slightly, from $30$ to $25$ in SIDM2v, and from $37$ to $26$ in SIDMx, but both remain significantly higher than the $5$ found in CDM.
Appendix E Candidates contributing to the GGSL cross sections
Figure 8: Bound subhalo mass ($M_{\rm sub}$) versus $r_{\rm max}$ for subhalos associated with the selected secondary lenses in the CDM (left), SIDM2v (middle), and SIDMx (right) simulations. Points are color-coded by the effective Einstein radius, defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ is the area enclosed by the corresponding critical curve. The subhalo mass range spans $10^{10}-10^{12}~{}\rm M_{\odot}/h$, and the resulting Einstein radii lie within $0.5^{\prime\prime}$$3^{\prime\prime}$, consistent with the findings of Ref. Dutra et al. (2025).
Ref. Dutra et al. (2025) demonstrates that the subhalos contributing to GGSL predominantly have bound masses in the range $10^{10}-10^{12}~{}\rm M_{\odot}/h$ and effective Einstein radii between $0.5^{\prime\prime}$ and $3^{\prime\prime}$, where the Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, with $A_{E}$ denoting the area enclosed by the critical curves. To show that the properties of the small-scale lenses in our study are consistent with these findings, we identify subhalos with effective $\theta_{E}>0.5^{\prime\prime}$ and plot their bound mass ($M_{\rm sub}$) against the radius at which their circular velocity peaks ($r_{\rm max}$) in Fig. 8, for the CDM (left), SIDM2v (middle), and SIDMx (right) simulations. The effective Einstein radii are indicated by the color scale. This result further supports the interpretation that SIDM models with mass segregation can produce subhalos consistent with GGSL observations.

### v2

@@ Appendix A SUPPLEMENTAL MATERIAL
Figure 7: Normalized density profiles for three $2\times 10^{10}~\rm M_{\odot}$ halos with median (blue), $1\sigma$ (green), and $2\sigma$ (orange) concentrations in the SIDM2c model, shown at increasing values of the gravothermal phase $\tau$. For comparison, the magenta curves show corresponding one-component SIDM profiles, and the initial NFW profile is shown in black. The red curves show the parametric model predictions for SIDM2c. The close match with the simulated curves demonstrates the quality of our calibration. The agreement among the three SIDM2c profiles (solid curves) in each panel demonstrates the conditioned universality of the model. Notably, the inner densities in the two-component case are consistently higher than in the one-component case apart from the late gravothermal phases ($\tau>1$), indicating that SIDM with mass segregation is more effective at enhancing the strong lensing signal.
Figure 8: Left: Normalized inner halo density evolution in SIDM2c for halos of $2\times 10^{10}~M_{\odot}$ and $10^{11}~M_{\odot}$ with various concentrations. The evolution time is normalized by the core collapse time calculated using Eq. 2 with $\sigma/m=14~\rm cm^{2}/g$. Middle: As in the left, but with the evolution time of all the simulations normalized to the first case in blue. Right: Model predictions (dash-dotted lines) versus simulation (shaded bands) for a $10^{11}~M_{\odot}$ halo ($1\sigma$ concentration, with baryons) at $t=0,3,6,9$ Gyr. Solid and dotted lines correspond to the halo and baryon profiles in the baryonic simulation, respectively. Dashed lines represent the halo profile in the DM-only simulation.
Appendix B A. A parametric model for the two-component model SIDM2c
B.1 Verification of the conditioned universality
In the main text, we illustrated the existence of a conditioned universality in SIDM models with mass segregation by showing the rescaled halo density profile at a normalized gravothermal phase of $\tau\equiv t/t_{c}=0.6$. Here, we further explore these universal features in Fig. 7 by showing density profiles at additional gravothermal phases: $\tau=0.1,0.2,0.4,0.8,1.0,$ and $1.08$. For simplicity, we only show profiles for the three halos of mass $2\times 10^{10}~\rm M_{\odot}$ and compare them with the one-component SIDM results (dashed magenta curves) at the same gravothermal phases. The one-component results are obtained using a parametric model calibrated to the conduction fluid model from Ref. Hou et al. (2025).
From Fig. 7, it is evident that the inner halo densities in the SIDM2c model are consistently higher than those in the one-component SIDM case across almost all gravothermal phases. Interestingly, in the late gravothermal phases ($\tau>1$), this difference diminishes, and the SIDM2c profiles become comparable to those of the one-component model. This result implies that the enhancement induced by SIDM2c does not stem from deeply core-collapsed halos, but rather from a broad population of halos across a range of gravothermal phases.
B.2 A parametric model for DM-only halos in SIDM2c
To make use of the conditioned universality, we construct a parametric model for the SIDM2c scenario based on the simulated $2\times 10^{10}~\rm M_{\odot}$ halo with $1\sigma$ concentration. We find the core collapse time in SIDM2c to be approximately that of a constant one-component SIDM model with $\sigma/m=14~\rm cm^{2}/g$. This allows us to compute the core collapse time as Balberg et al. (2002); Essig et al. (2019):
$t_{c}$ $=$ $\frac{150}{C}\frac{1}{(\sigma/m)\rho_{s}r_{s}}\frac{1}{\sqrt{4\pi G\rho_{s}}},$ (2)
where we adopt $C=0.75$, and $\rho_{s}$ and $r_{s}$ are the NFW scale parameters for the CDM halo under consideration.
Figure 8 (left) illustrates the evolution of the inner halo density, where time is normalized by the core collapse time $t_{c}$ defined in Eq. 2. For each simulation, the inner density is computed at $r=0.05r_{s}$ using the NFW scale parameters $\rho_{s}$ and $r_{s}$, and normalized by $\rho_{s}$. Without the time rescale by $t_{c}$, the core collapse times for these halos range from 15 to 110 Gyr. A clear, unified evolutionary trend emerges after normalization, despite a roughly 30% uncertainty in $t_{c}$, which is comparable to the uncertainty in the one-component case.
Note that the conditioned universality in theory is independent of the $t_{c}$ formula. In Fig. 8 (middle), we rescale the evolution time in all the cases to the core collapse time of the first case shown. The agreement among the curves significantly improves over the left panel, hence future work can involve calibrating a more accurate expression for $t_{c}$ to reduce uncertainty in the parametric model.
The parametric model is based on the density profile introduced in Ref. Hou et al. (2025), calibrated by fitting the evolution of its profile parameters to the simulated SIDM2c halo.
The model is parameterized as
$\rho_{\rm PSIDM2c}(r|\rho_{s,0},r_{s,0},\tau)$ (3)
where $\rho_{s}=\rho_{s}(\rho_{s,0},\tau)$, $r_{s}=r_{s}(r_{s,0},\tau)$, $r_{c}=r_{c}(r_{s,0},\tau)$, $\beta=\beta(\tau)$ and $\gamma=\gamma(\tau)$ are functions of the phase parameter $\tau\in(0,1.08)$. The profile parameters as functions of $\tau$ and are fitted to be
$\rho_{s}(\tau)/\rho_{s,0}$ $=$ $\exp\left(-15.74\ln(3.748\tau+1)-0.4013+31.26\tau-15.58\tau^{2}+2.999\tau^{3}\right),$ (4)
$r_{s}(\tau)/r_{s,0}$ $=$ $1.116+7.066\tau+38.89\tau^{2}-217.3\tau^{3}+484.3\tau^{4}-557.2\tau^{5}+328.1\tau^{6}-77.84\tau^{7},$ (5)
$r_{c}(\tau)/r_{s,0}$ $=$ $0.5878\sqrt{\tau}-0.6073\tau+1.102\tau^{3/2}-1.384\tau^{2}-1.492\tau^{3}+5.162\tau^{4}-4.951\tau^{5}+1.634\tau^{6},$ (6)
$\beta(\tau)$ $=$ $1.128+9.554\tau-30.26\tau^{2}+124.9\tau^{3}-109.4\tau^{4}-181.9\tau^{5}+276.5\tau^{6},$ (7)
$\gamma(\tau)$ $=$ $0.2651+3.941\sqrt{\tau}+0.2936\tau-9.062\tau^{2}+14.35\tau^{3}-10.22\tau^{4}+2.785\tau^{5}$ (9)
$\quad+\frac{1}{\ln(0.001)}\left[1-0.2651\right]\ln(\tau+0.001).$
Based on the parameterized density profiles, we also extract the evolution of $V_{\rm max}$ and $R_{\rm max}$, where $V_{\rm max}$ denotes the maximum rotation velocity and $R_{\rm max}$ its corresponding radius. We obtain the following fitting functions to describe their evolution:
$V_{\rm max,SIDM2c}(\tau)/V_{\rm max,0}$ $=$ $1-0.2850\,\tau+0.5936\,\tau^{2}+0.01723\,\ln(221.2\,\tau+1)-0.3175\,\tau^{3}+0.2157\,\tau^{4},$ (10)
$R_{\rm max,SIDM2c}(\tau)/R_{\rm max,0}$ $=$ $1+[1-(1+e^{-20(\tau-0.3)})^{-1}]P_{1}(\tau)+(1+e^{-20(\tau-0.3)})^{-1}P_{2}(\tau)$ (11)
$P_{1}(\tau)$ $=$ $-2.305\,\tau+17.23\,\tau^{2}-45.64\,\tau^{3}$ (12)
$P_{2}(\tau)$ $=$ $-0.7538\,\tau-2.228\,\tau^{2}+3.508\,\tau^{3}-1.429\,\tau^{4}.$ (13)
To map CDM halos from cosmological simulations to their SIDM2c counterparts, we incorporate the effect of accretion histories using the integral approach proposed in Ref. Yang et al. (2024). In this approach, the gravothermal evolution is integrated along the evolutionary trajectories of $V_{\rm max,CDM}$ and $R_{\rm max,CDM}$ in the CDM simulation, yielding their predicted values in SIDM2c as
$V_{\rm max}(t)$ $=$ $V_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dV_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{c}(t^{\prime})}\frac{dV_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}}$
$R_{\rm max}(t)$ $=$ $R_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dR_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{c}(t^{\prime})}\frac{dR_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}},$ (14)
where $t_{f}$ is the halo formation time, chosen right after the halo’s exponential accretion phase. The derivatives $dV_{\rm max,SIDM2c}(\tau)/d\tau$ and $dR_{\rm max,SIDM2c}(\tau)/d\tau$ are obtained by differentiating the corresponding functions in Eq. 10, with the initial values $V_{\rm max,0}$ and $R_{\rm max,0}$ replaced by $V_{\rm max,CDM}(t^{\prime})$ and $R_{\rm max,CDM}(t^{\prime})$, respectively. After obtaining $V_{\rm max}(t)$ and $R_{\rm max}(t)$ for the corresponding $\tau$, the full SIDM2c density profile can be constructed by solving for the NFW parameters in Eq. 4. Further technical details are available in Ref. Yang et al. (2025b), which validates this approach using a population of halos from cosmological simulations. This method has also been applied to the SIDM Concerto simulation suite for tests across multiple mass scales Nadler et al. (2025).
B.3 Incorporating the influence of baryons
As baryons radiate their energy away from their host halos, they are generically more compact. The presence of a deeper baryon potential contracts dark matter and can boost the halo’s dynamical evolution.
To incorporate the effect of baryons on the halo density profile, we adopt a simplified approach by calibrating to the two simulated halos with baryons. Both simulations have halo mass $10^{11}~\rm M_{\odot}$ and the stellar content is set by median stellar-to-halo mass and size-mass relations Moster et al. (2013); Carleton et al. (2019).
Our approach is parametric, utilizing the form factor formula from Ref. Yang (2024), defined as
${\cal F}$ $=$ $\left(\frac{1}{\hat{r}_{\rm eff}}+\frac{20\hat{\rho}_{H}\hat{r}_{H}^{3}}{\hat{r}_{\rm eff}(\hat{r}_{\rm eff}+\hat{r}_{H})^{2}}\right)^{-1}\left(1+1.6\frac{\hat{\rho}_{H}\hat{r}_{H}^{2}}{2}\right)^{-\frac{1}{2}},$
where $\rho_{s}$, and $r_{s}$ are the NFW scale parameters and $\rho_{H}$, and $r_{H}$ are the Hernquist scale parameters. The dimensionless quantities are defined as $\hat{\rho}_{H}=\rho_{H}/\rho_{s}$, $\hat{r}_{H}=r_{H}/r_{s}$, and $\hat{r}_{\rm eff}=r_{\rm eff}/r_{s}=(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{3}/2)/(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{2}/2)$. To obtain predictions that incorporate baryonic effects, we simply multiply this factor by the core size $r_{c}(\tau)$ in the parametric model for DM-only halos. While schematic, this method accounts for the baryon-induced contraction effect, which scales with the stellar-to-halo mass ratio, and smoothly reduces to the no-baryon limit as $\tau$ approaches zero.
In the main text, we have illustrate the simulation results for a median concentration halo in Fig. 3 middle panel, where baryons have growing cores along with the dark matter. The parametric model predictions are tested against the simulation results at different times. Here we provide another example for a more compact halo. In the right panel of Fig. 8, we compare the parametric model predictions with the simulation results for a $M^{11}~\rm M_{\odot}$ halo ($1\sigma$ concentration). The profiles, shown in progressively darker colors, correspond to $t=0,3,6,$ and $9$ Gyr, respectively. While small deviations in shape are present, the model effectively captures the growth of inner densities in the baryonic case (solid lines) compared to the DM-only cases (dashed lines). The baryon profiles themselves (dotted lines) remain largely unchanged in this benchmark.
Appendix C B. Calculation of strong lensing effects
Figure 9: Density profiles of cluster subhalos with $M>10^{11}\,M_{\odot}/h$ in CDM, SIDM2c, and SIDM2c+baryons from parametric model predictions (dashed). Solid curves show corresponding fits using the PSIDM model of Refs. Hou et al. (2025). Subpanels display the relative deviation between simulation and fit, defined as $2\,(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The parametric model accurately captures the density structure of simulated subhalos across a broad mass range.
After obtaining density profiles from either the SIDM2c parametric model or from simulations, we compute the corresponding lensing signatures. To efficiently obtain lensing predictions for a population of halos, we use the parametric lensing model for SIDM halos developed in Ref. Hou et al. (2025), except for the host halo. Although originally developed for one-component SIDM halos, this model can be extended to more general density profiles by treating it as a three-parameter fitting template ($\rho_{s,0}$, $r_{s,0}$, and $\tau$). Notably, the gravothermal evolution involves halo inner densities that vary from $r^{0}$ to $r^{-2.5}$, providing considerable modeling flexibility. A superposition of two such profiles is capable of fitting the more complex structures that arise in the inner halo region.
Specifically, the parametric SIDM (PSIDM) density profile introduced in Ref. Hou et al. (2025) has the same parametric form as Eq. 3. The evolution of its halo parameters is governed by the functions given in that reference. For CDM halos, a single PSIDM profile is sufficient for an accurate fit, as it reduces to the NFW profile in the limit $\tau\to 0$. For SIDM2c halos, we use a superposition of two PSIDM profiles to achieve an accurate representation.
Figure 9 shows the fitting performance for cluster subhalos in three models: from left to right, CDM (blue), SIDM2c (red), and SIDM2c+baryons (magenta). Only halos with masses above $10^{11}~\rm M_{\odot}$ are shown. The lower panels illustrate the relative difference between the simulated (Sim) and fitted (Fit) profiles, computed as $2(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. Clearly, the two PSIDM model reproduces the simulated halo profiles with high accuracy, even in the presence of baryonic effects.
Based on the fitted profiles, Ref. Hou et al. (2025) provides analytic expressions for the surface density, lensing potential, and deflection angles of each halo. The linearity of the lensing equations allows us to superimpose their contributions for more comprehensive lensing predictions.
The cluster host halos are modeled differently from the other halos. While the internal ellipticity of other halos contributes subdominantly, the mass and ellipticity of the host halo itself significantly affect the overall lensing signal and must be incorporated. We model this effect by applying a coordinate transformation that introduces a fixed ellipticity into the projected mass distribution.
The ellipticity of each host halo is measured by projecting the associated particles onto the lens plane, taken here as the xxyy plane of the simulation box (in the fiducial case without rotation). We define the axis ratio $q=b/a$ of the projected mass distribution and compute the elliptical radial coordinate as Kormann et al. (1994)
$s=\sqrt{q^{2}(x\cos\phi+y\sin\phi)^{2}+(-x\sin\phi+y\cos\phi)^{2}},$
where $\phi$ is the angle between the $x$-axis and the semi-major axis of the lens. Substituting $s$ for $r$ in the circular lens model yields an elliptical lens, preserving the enclosed mass within isodensity contours.
We apply a fast Fourier transform to the elliptical surface density to compute the deflection angle maps for the host halos. For the CDM, SIDM2v, and SIDMx clusters, the measured axis ratios are $q=0.39$, $0.42$, and $0.47$, respectively. These large ellipticities enhance the GGSL effects associated with the host halos.
Appendix D C. Cluster density profiles and constraint
Figure 10: Left: Density profiles of cluster host halos in the CDM, SIDM2v, and SIDMx simulations (dashed). The profiles are fitted by two PSIDM density profiles (solid) Hou et al. (2025). Right: Comparison of cluster density profiles at $t=5~\rm Gyr$ for the two-component SIDM2v model (solid) and the one-component SIDM01 model (dashed). The initial CDM halo (black) and baryons (orange) are also shown. The inner densities for SIDM01 are consistently shallower than those for SIDM2v, both with and without baryons, indicating that SIDM2v remains consistent with cluster-scale constraints.
@@ B.3 Incorporating the influence of baryons
Figure 10 (left) presents the density profiles of host halos from cosmological simulations. As a result of the merging environment and its interplay with inter- and intra-species scattering, we observe intriguing twists in the density profiles of both SIDM2v and SIDMx, which complicates the assessment of cluster-scale constraints.
To obtain a robust assessment, we perform isolated simulations for the cluster halo under SIDM2v, using the scale parameters of the CDM cluster halo as a reference. We consider two contrasting scenarios with and without baryons. The baryonic component, with a mass of $7.47\times 10^{11}~\rm M_{\odot}$, is modeled as a Hernquist profile with $r_{H}=34.1~\rm kpc$. These parameters align with the median stellar-to-halo mass relation and median size-mass relation Moster et al. (2013); Carleton et al. (2019). We use $5\times 10^{6}$ dark matter simulation particles in both cases to better resolve the inner halo regions. Baryons are treated as collisionless stars with particle masses equal to those of the dark matter.
Figure 10 (right) shows the simulation results (solid) at $t=5~\rm Gyr$, compared with the one-component density profiles (dashed) from the parametric models of Ref. Yang (2024) with bayrons and Ref. Yang et al. (2024) without baryons. The resulting density profiles in SIDM2v are denser than those of the one-component SIDM01 model ($\sigma/m=0.1~\rm cm^{2}/g$), even in the presence of baryons. Since observational constraints depend on the inner mass distribution, our results imply that clusters in SIDM2v are less constrained than those in SIDM01. In both cases, baryons induce a small contraction that does not alter this conclusion.
Appendix E D. Uncertainties in the GGSL cross section
Our simulations and analyses are subject to several systematic and statistical uncertainties. In this section, we quantify the dominant uncertainty by analyzing the lens system from different viewing angles, discuss limitations of our numerical methods, and demonstrate the robustness of our primary findings.
To isolate the signal of the relative enhancement from SIDM with mass segregation compared to CDM, we eliminate the inherent randomness in hierarchical structure formation by using identical initial conditions for all simulations. This approach ensures a one-to-one mapping for massive halos across the CDM and SIDM runs, provided SIDM does not substantially alter their accretion histories. Therefore, the relative enhancement can be quantified on a halo-by-halo basis. This design intentionally removes halo-to-halo variance, which would otherwise obscure observable signatures. Details such as variations between different cluster halos or the specific spatial distribution of subhalos are rendered marginal in our analysis.
The dominant uncertainty in our GGSL cross section computation arises from analyzing the lens system from different lines of sight. This is significant because strong lensing is sensitive to shear, and the projected ellipticity of our lens changes considerably with viewing angle. In our fiducial setup, the lens lies in the xxyy plane and is observed along the $z$-axis. We then rotate the viewing angle about the $x$- and $y$-axes in increments of $18^{\circ}$, up to $90^{\circ}$, generating 11 distinct realizations for each cosmological simulation. The statistics of the lensing effects are derived from the ensemble of these realizations. We summarize these results in Table 2 of the main text.
We further quantify the statistical differences in GGSL cross sections among the models. This is done using Welch’s unequal variances $t$-test, which is suitable for comparing the 11 independent realizations per model that may have different variances. The $p$-values and corresponding Gaussian significance in unit of $\sigma$ from this test are presented in Table 3.
The comparison between CDM and Model-2v yields a $2.14\sigma$ difference. All other model comparisons with the CDM baseline show highly significant differences, exceeding $4\sigma$. These results provide strong evidence that our proposed mechanism—SIDM with mass segregation—can produce a significant excess of GGSL events precisely in the regime of observational interest.
Table 3: Statistical significance of GGSL cross section comparisons.
Comparison $p$-value significance ($\sigma$)
CDM vs Model-2v 0.033 2.14
CDM vs Model-x $3.78\times 10^{-6}$ 4.62
CDM vs Model-2v+baryons $1.72\times 10^{-5}$ 4.30
CDM vs Model-x+baryons $1.03\times 10^{-5}$ 4.41
Figure 11: Histograms for the Einstein radii of the contributing secondary critical lines in CDM (blue), Model-2v (red), and Model-x (magenta) models. The corresponding cases that incorporate baryonic effects are shown as dashed lines. A vertical dotted line indicates the veto cut $\theta_{\rm cut}=3^{\prime\prime}$, which is applied to isolate GGSL effects in regions containing observed critical lines. The solid black line shows the observed number density, as taken from Fig. 7 of Ref. Meneghetti et al. (2022).
Figure 12: Convergence test for a $10^{11}~\rm M_{\odot}$ halo with $1\sigma$ concentration (top) and a $10^{13}~\rm M_{\odot}$ halo with median concentration (bottom). For each halo, we consider $N=5\times 10^{4}$ (blue), $10^{5}$ (orange), $5\times 10^{5}$ (red), and $10^{6}$ (magenta) simulation particles. The figure shows the evolution of their inner densities at several enclosed radii.
Appendix F E. Candidates contributing to the GGSL cross sections
Refs. Meneghetti et al. (2022, 2023); Dutra et al. (2025) demonstrate that the subhalos contributing to GGSL predominantly have bound masses in the range $10^{10}-10^{12}~\rm M_{\odot}/h$ and effective Einstein radii $\theta_{E}$ between $0.5^{\prime\prime}$ and $3^{\prime\prime}$. Here, the Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ is the area enclosed by the critical curve.
To confirm that the small-scale lenses in our study are consistent with these findings, we identify subhalos with $\theta_{E}>0.5^{\prime\prime}$ and plot their $\theta_{E}$ distribution in Fig. 11 for the models in Table 2. The number density is obtained by dividing the total number of contributing critical lines by the number of considered systems (11) and by the fiducial calculation area of $100^{\prime\prime}\times 100^{\prime\prime}$. While a few systems have large Einstein radii ($\theta_{E}>3^{\prime\prime}$), the majority of candidates lie in the range $\theta_{E}<3^{\prime\prime}$, in good agreement with observations. Notably, the number densities in our benchmark models, especially those including baryons, show distributions that match the observational data. This agreement further supports the interpretation that SIDM models with mass segregation can produce subhalo populations consistent with GGSL observations.
The candidates with $\theta_{E}>3^{\prime\prime}$ correspond to massive systems, typically with masses exceeding $10^{13}~\rm M_{\odot}$. A stronger core formation driven by intra-species interaction or AGN feedback could produce shallower inner density profiles, shrinking their Einstein radii. Since the physics governing these higher-mass systems operates at a different scale, their properties can be adjusted naturally through either dark matter or baryonic physics. We defer a fully consistent treatment of these systems to future work.
Appendix G F. Convergence test for halos in the cosmological simulation
This work involves two types of simulations. The controlled simulations are high-resolution, typically using $10^{6}$ or more particles per halo. The cosmological simulations, on the other hand, have a fixed average particle mass of $m_{0}\equiv(m_{H}+m_{L})/2=10^{8}~\rm M_{\odot}$. This resolution implies that a $5\times 10^{12}~\rm M_{\odot}$ halo is resolved with only $5\times 10^{4}$ particles.
To demonstrate that our cosmological simulation program can adequately resolve halos down to this mass scale ($5\times 10^{12}~\rm M_{\odot}$), we performed a series of controlled convergence tests. We simulated a $10^{11}~\rm M_{\odot}$ halo (with $1\sigma$ concentration) and a $10^{13}~\rm M_{\odot}$ halo (with median concentration) at different resolutions: $N=5\times 10^{4}$, $10^{5}$, $5\times 10^{5}$, and $10^{6}$ particles.
Figure 12 presents the results. It shows the mean inner density evolution at $1$, 22, and $3$ kpc for the $10^{11}~\rm M_{\odot}$ halo, and at $4$, 88, and $12$ kpc for the $10^{13}~\rm M_{\odot}$ halo. The tracks for all resolutions converge with the highest-resolution ($N=10^{6}$) result. While simulations with fewer particles show increased fluctuations, their overall evolution trend aligns with the high-resolution result.
Appendix H G. Summary of controlled simulations
We summarize in Table 4 the parameters of the high-resolution controlled simulations in this work.
Simulation $M_{h}({\rm M_{\odot}})$ $\rho_{s}({\rm M_{\odot}/kpc^{3}})$ $r_{s}({\rm kpc})$ $M_{s}({\rm M_{\odot}})$ $r_{H}({\rm kpc})$ SIDM model $N_{\rm part}$
M02 $2\times 10^{10}$ $1.03\times 10^{7}$ $4.53$ 0 0 SIDM2c $10^{6}$
M02p1 $2\times 10^{10}$ $1.94\times 10^{7}$ $3.51$ 0 0 SIDM2c $10^{6}$
M02p2 $2\times 10^{10}$ $3.71\times 10^{7}$ $2.73$ 0 0 SIDM2c $10^{6}$
M05 $5\times 10^{10}$ $4.05\times 10^{7}$ $3.58$ 0 0 SIDM2c $10^{6}$
M08 $8\times 10^{10}$ $3.58\times 10^{7}$ $4.39$ 0 0 SIDM2c $10^{6}$
M11 $10^{11}$ $6.89\times 10^{6}$ $9.10$ 0 0 SIDM2c $10^{6}$
M11 $10^{11}$ $6.89\times 10^{6}$ $9.10$ 0 0 SIDM03 $10^{6}$
M11p1 $10^{11}$ $1.29\times 10^{7}$ $7.07$ 0 0 SIDM2c $10^{6}$
M11m1 $10^{11}$ $3.73\times 10^{6}$ $11.7$ 0 0 SIDM2c $10^{6}$
M11B $10^{11}$ $6.89\times 10^{6}$ $9.10$ $10^{9}$ 0.768 SIDM2c $10^{6}+8281$
M11p1B $10^{11}$ $1.29\times 10^{7}$ $7.07$ $10^{9}$ 0.768 SIDM2c $10^{6}+8420$
M15 $1.79\times 10^{15}$ $3.09\times 10^{5}$ $860$ 0 0 SIDM2v $5\times 10^{6}$
M15B $1.79\times 10^{15}$ $3.09\times 10^{5}$ $860$ $7.47\times 10^{11}$ $34.1$ SIDM2v $5\times 10^{6}+1556$
Table 4: Parameters of the controlled N-body simulation halos. The labels “p1”, “p2”, and “m1” in the benchmark names refer to simulations with concentrations of $1\sigma$, $2\sigma$, and $1\sigma$ relative to the median. The labels “B” imply the presence of baryons. Halo masses are determined as the mass enclosed within radii where the density reaches 200 times the critical density ($R_{200}$), based on the scale density ($\rho_{s}$) and scale radius ($r_{s}$) of the NFW profile. The total halo masses in the simulations are slightly higher due to the smooth truncation of the density profiles. Stellar masses are calculated as total masses using the scale density ($\rho_{H}$) and scale radius ($r_{H}$) of the Hernquist profile. Dark matter and stellar simulation particles have equal mass. $N_{\rm part}$ denotes a summation of their numbers.